Having trouble with infinite series... Actually with all algebra. But,
Question: Find the sum of these infinite series. If it is not possible, write "no sum" and explain why.
1) 1024 + 512 + 256 +128 + ...
2) 1/9 + 1/3 + 1 + 3 +...
3) 16 + 4 + 1 + 1/4 +...

Answer by richwmiller(17219)

An infinite geometric series converges if its common ratio r satisfies –1 < r < 1.
Sum=(a1/1-r)
1)r=1/2
Sum=(1024/1-1/2=1024/1/2=2048
2) no sum r=3 which is greater than 1
3) r=1/4
Sum=16/(1-1/4)=16/(3/4)=64/3=21.333